Can you please solve the following Linear Equations for the unknown:
3𝑧 + 2 = 5𝑧 − 2
𝑑 – 3 = 2𝑑 + 6

Answer by Cromlix(4381)

Hi there,
3z + 2 = 5z - 2
3z - 5z = -2 - 2
-2z = -4
z = -4/-2
z = 2
........
d - 3 = 2d + 6
d - 2d = 6 + 3
-d = 9
Multiply both sides by -1
d = -9
Hope this helps :-)